# Yuqin Dai1,3,    Guoqing Wang3∗,    Yuan Wang2,3∗,    Kairan Dou4,5,    Kaichen Zhou4 Zhanwei Zhang2,3 ,    Shuo Yang6,    Fei Tang2,3,    Jun Yin1,    Pengyu Zeng1,    Zhenzhe Ying3

### v1

@@ Abstract
Large Language Models (LLMs) empowered with retrieval mechanisms have achieved strong progress in open-domain question answering (QA). Yet, the conventional retrieve-then-answer paradigm often suffers from two key limitations: (1) low signal-to-noise ratio in retrieved evidence, where useful information is buried under irrelevant content, and (2) error accumulation in multi-hop reasoning when incomplete or noisy passages are involved. To address these challenges, we present EviNote-RAG, an agentic RAG framework that introduces a structured retrieve–note–answer pipeline. Instead of directly reasoning over raw retrievals, the model is trained to compose Supportive-Evidence Notes (SENs), concise, human-like notes that preserve only answer-relevant information, highlight uncertainty, and explicitly state when no useful evidence exists. This distillation process is further reinforced by the Evidence Quality Reward (EQR), an entailment-based signal that evaluates whether SENs logically support the final answer. Together, SENs and EQR guide the model toward faithful and robust reasoning, while reducing the impact of noise. Experiments on in-domain and out-of-domain QA benchmarks show that EviNote-RAG consistently outperforms strong baselines in accuracy, generalization, and training stability. In particular, it achieves state-of-the-art results while enhancing robustness and efficiency, yielding relative F1 gains of 20% on HotpotQA (+0.093), 40% on Bamboogle (+0.151), and 91% on 2Wiki (+0.256) via denser rewards and reduced verbosity.
@@ 1 Introduction
Large Language Models (LLMs) have evolved from next-token predictors into general-purpose systems capable of advanced reasoning, code generation, and open-domain question answering (Brown et al., 2020; Chowdhery et al., 2022; Verma et al., 2024b; Zheng et al., 2025; Jiang et al., 2025). Since the factual knowledge of LLMs is fixed during pre-training, they are prone to generating incorrect or outdated information when deployed in real-world tasks where knowledge evolves rapidly (Ji et al., 2022; Zhang et al., 2025). To address this limitation, Retrieval-Augmented Generation (RAG) methods have emerged by incorporating a search component that supplies up-to-date external evidence during both training and inference, enabling models to ground their responses in timely information and improve factual consistency.
Despite recent advances, most retrieval-augmented generation (RAG) systems—from early frameworks (Lewis et al., 2020; Guu et al., 2020; Jiang et al., 2023b; Guan et al., 2025) to recent RL-based models (Zhang et al., 2024a; Wei et al., 2025a; Song et al., 2025b; Jin et al., 2025; Li et al., 2025a; Kaelbling et al., 1996)—still adopt a retrieve-then-answer paradigm, deriving conclusions directly from raw, often noisy passages. This design gives rise to two persistent challenges: (1) Low Signal-to-Noise Ratio (SNR). Retrieved evidence often includes substantial irrelevant content, making answer-supporting information sparse and difficult to identify (Shi et al., 2023; Jin et al., 2024). (2) Error Accumulation. Reasoning errors compound when inference is required across incomplete or noisy evidence, a problem especially severe in multi-hop QA (Shi et al., 2023). To mitigate these issues, prior works have explored several directions, such as enhancing retrieval quality (Gutiérrez et al., 2025; Shao et al., 2025; Sun et al., 2025a; Wang et al., 2025), filtering or compressing retrieved content through re-ranking, summarization, or external models (Glass et al., 2022; Dong et al., 2024; Asai et al., 2022; Xu et al., 2023; Kim & Lee, 2024; Laban et al., 2024; Kim et al., 2024; Jiang et al., 2023a; Zhang et al., 2024b), and applying supervised fine-tuning (SFT) to improve information extraction over specific corpora (Liu et al., 2024; Zhang et al., 2024a). However, while these approaches alleviate noise and reduce reasoning burden to some extent, they remain constrained by dataset dependence or reliance on external heuristics. Fundamentally, the model itself still lacks the ability to reliably discern and selectively exploit useful information.
@@
To address the core limitation of existing RAG systems, we propose EviNote-RAG, an end-to-end agentic RAG framework that restructures the pipeline into a retrieve–note–answer process (Fig. 1). EviNote-RAG trains LLMs to generate Supportive-Evidence Notes (SENs), concise abstractions that preserve only answer-critical information, discard irrelevant content, and explicitly indicate when no relevant evidence is found. Each SEN further highlights key and uncertain information, echoing human note-taking strategies to improve focus and reduce misleading reasoning. Most importantly, we formulate evidence selection as a reinforcement learning problem: through the Evidence Quality Reward (EQR), a lightweight entailment judge evaluates how well each SEN supports the final answer. This reward signal encourages the model to freely explore strategies for evidence extraction while being guided toward more accurate and faithful use of information. As a result, EviNote-RAG achieves end-to-end optimization that reduces noise, mitigates error accumulation, and enables the model to learn an effective strategy for precise information utilization.
@@ 2 Related Work
2.1 Instruction-Guided Agentic RAG
Instruction-guided methods (Yao et al., 2023; Jeong et al., 2024; Amplayo et al., 2022) enhance RAG by designing prompts that automate retrieval and guide multi-step reasoning (Wu et al., 2025; Song et al., 2025b). These approaches (Zhang et al., 2024a; Jiang et al., 2023b) typically decompose questions into sub-problems, retrieve external knowledge, and synthesize structured answers (Asai et al., 2024; Lewis et al., 2020; Guu et al., 2020). Other works (Press et al., 2022; Trivedi et al., 2022a) integrate retrieval directly into the reasoning loop, while more recent efforts (Li et al., 2025a; Alzubi et al., 2025; Feng et al., 2025; Yue et al., 2024; Verma et al., 2024a; Guan et al., 2025) interleave retrieval and reasoning adaptively. Complementary test-time filtering strategies include passage re-ranking (Glass et al., 2022; Dong et al., 2024), external-model filtering (Asai et al., 2022; Xu et al., 2023; Kim & Lee, 2024; Laban et al., 2024), instruction-guided summarization (Kim et al., 2024; Jiang et al., 2023a; Zhang et al., 2024b), and stronger retrievers (Gutiérrez et al., 2025; Shao et al., 2025; Sun et al., 2025a; Wang et al., 2025). Despite these advances, prompt-based approaches inherently depend on the foundation model’s generalization ability in RAG, which remains limited. Our framework instead employs a post-training, reward-driven objective that explicitly shapes information selection and reasoning, leading to more faithful and task-adapted answers.
@@ 2.2 Reward-Guided Agentic RAG
Reward-guided approaches (Zhang et al., 2024a; Huang et al., 2025; Wei et al., 2025a) employ reinforcement learning (Kaelbling et al., 1996) to optimize reasoning policies through scalar feedback derived from task performance. Early work (Nakano et al., 2021) demonstrated that reward signals can effectively guide multi-step retrieval and improve factual accuracy. Subsequent works (Qi et al., 2025; Chen et al., 2025; Jin et al., 2025) broadened this paradigm by exploring diverse scenarios and tool-use (Wei et al., 2025b; Dai et al., 2025), further underscoring the performance gains enabled by reward-based supervision for retrieval-augmented reasoning. However, most existing reward-guided methods (Wu et al., 2025; Song et al., 2025b; Dai et al., 2025; Jin et al., 2025) operate directly on raw, often noisy passages, which leads to low signal-to-noise ratio (Shi et al., 2023; Jin et al., 2024) and error accumulation across multi-hop reasoning. Some approaches (Liu et al., 2024; Zhang et al., 2024a) attempt to alleviate this via supervised fine-tuning (SFT) on specific corpora, but such gains are limited by dataset dependence. EviNote-RAG tackles this limitation by using supportive-evidence notes to structure retrieved information and by applying supervision that enforces logical consistency between the notes and the final answers. These mechanisms together promote more reliable reasoning and improve answer accuracy.
@@ 3 Methodology
3.1 Preliminary
In each step, the agent decides whether to retrieve information or to stop and generate the final answer. This multi-turn process is an agentic Retrieval-Augmented Generation (RAG) task, which can be modeled as a sequential decision-making problem with a Markov Decision Process (MDP).
Formally, given a user instruction $I$, we define this MDP as a tuple $(\mathcal{S},\mathcal{A},\mathcal{R},\mathcal{P})$, where $\mathcal{S}$ denotes the set of states, $\mathcal{A}$ is the set of actions, $\mathcal{R}$ is a scalar reward function, and $\mathcal{P}(s^{\prime}\mid s,a)$ defines the probability of transitioning to state $s^{\prime}$ given the current state $s$ and action $a$. Specifically, at each timestep $t$, the state $s_{t}\in\mathcal{S}$ includes two parts: (1) the text generated so far, and (2) any external information the agent has retrieved up to that point. Given state $s_{t}$, the agent selects an action $a_{t}\in\mathcal{A}$ according to its policy $\pi_{\theta}$, parameterized by $\theta$:
$a_{t}=\pi_{\theta}(s_{t}).$ (1)
We explicitly introduce a special search action $a^{\prime}\in\mathcal{A}$, which triggers retrieval from an external search engine. When this action is selected, the environment provides retrieved content $d_{t}$, which is then appended to the current state. The state transition is thus defined as follows:
$s_{t+1}=\begin{cases}\operatorname{concat}(s_{t};a_{t},d_{t}),&\text{if }a_{t}=a^{\prime},\\ \operatorname{concat}(s_{t};a_{t}),&\text{otherwise},\end{cases}$ (2)
where $\operatorname{concat}$ denotes sequence concatenation. At each timestep $t$, the agent receives a reward $r_{t}=\mathcal{R}(s_{t},a_{t})$ from the environment, based on the current state $s_{t}$ and the selected action $a_{t}$.
$r_{t}=\mathcal{R}(s_{t},a_{t}),$ (3)
The reward function $\mathcal{R}$ will be detailed in Sec. 3.3. Finally, an episode terminates either when the task is successfully completed or when the maximum number of retrieval steps is reached.
@@ 3.2 EviNote-RAG Pipeline
To address the challenge of low information utilization, EviNote-RAG introduces a dedicated note-taking phase that enables the model to identify and summarize supportive evidence. As illustrated in Fig. 2, the model produces Supportive-Evidence Notes (SENs) that preserve only the information necessary to answer the question. To ensure that each SEN genuinely supports the final answer, we incorporate an Evidence Quality Reward (EQR) based on entailment supervision, which guides the model toward faithful and focused content selection. This mechanism leads to more accurate answers. The complete method is detailed in the following section.
Paradigm of EviNote-RAG Inference.
Our EviNote-RAG proceeds by first having the agent reason through the task, enclosed within <think> tags (Guo et al., 2025). Based on this reasoning, the agent may issue queries to a search engine using <search> tags. Retrieved results are returned by the environment and encapsulated within <information> tags, which are appended to the prompt for the next turn. It is worth noting that the retrieved information often has a low signal-to-noise ratio (SNR), containing substantial distracting noise that may mislead the model into incorrect reasoning. To improve information utilization, our method generates Supportive-Evidence Notes (SENs) enclosed in <summary> tags. These SENs act as a filtering mechanism, discarding irrelevant content and retaining only the supportive evidence necessary to derive the correct answer, thereby enhancing both information utilization and reasoning accuracy. Finally, once sufficient evidence is collected, the agent concludes by generating a final answer enclosed in <answer> tags.
Generating Supportive-Evidence Notes (SENs)
To enhance information use, SENs include two types of annotations: key information (marked with *) and uncertain information (marked with –). These annotations mirror human strategies for emphasizing importance and signaling ambiguity, helping prevent misunderstanding and improving reasoning clarity. To guide the model in producing high-quality SENs, we introduce the Evidence Quality Reward (EQR), which provides entailment-based feedback to help the model focus on answer-relevant content. We describe EQR in detail in the following section.
3.3 Training Strategy
Evidence Quality Reward (EQR)
@@ 3.3 Training Strategy
To encourage the generation of high-quality Supportive-Evidence Notes (SEN), we introduce an Entailment Judge as a source of supervision. The underlying intuition is straightforward: a well-formed SEN should provide sufficient grounds for logically inferring the ground-truth answer. We realize the Entailment Judge through a lightweight Natural Language Inference (NLI) model (e.g., DistilBERT (Sanh et al., 2019)), which evaluates whether the final SEN entails the correct answer. To be more specific, we first construct an answer claim $h$ that asserts the ground-truth answer $\text{ANS}_{\text{gt}}$ is the correct answer to the question $q$. We then use the final SEN $\text{SEN}_{\text{last}}$ as the input text, and evaluate whether it logically supports $h$ using the Entailment Judge model $\mathcal{M}_{\text{Judge}}$, as shown below:
@@
  • ✓ indicates that a condition is satisfied, and ✗ indicates violation;
  • format is valid if the output includes an explicit <answer> tag, contains at least one <summary> tag, and follows the prescribed tagging schema;
  • answer is correct if it exactly matches the ground-truth answer (Exact Match);
  • note-taking is satisfied if the model annotates the SENs with “*” for key information and “-” for uncertain information, as described in our SEN design.
This reward design ensures that the model is gently guided to perform structured note-taking when its QA capability is still developing (through a small reward of $0.1+R_{\text{EQR}}$), and increasingly incentivized to generate precise, logically supported answers when it becomes more reliable (through $1+R_{\text{EQR}}$). Note that this strategy integrates the Evidence Quality Reward (EQR) $R_{\text{EQR}}$, which provides entailment-based feedback to further emphasize relevance and faithfulness in the final generated answers.
@@ 4.1 Datasets
We evaluated EviNote-RAG on seven widely used Question Answering (QA) benchmark datasets: (1) In-Domain Datasets consist of NQ (Kwiatkowski et al., 2019) and HotpotQA (Yang et al., 2018). These datasets are considered in-domain because they originate from the same question distribution, allowing for a direct comparison on familiar tasks. (2) Out-of-Domain Datasets: Out-of-Domain Datasets include PopQA (Mallen et al., 2022), TriviaQA (Joshi et al., 2017), 2WikiMultiHopQA (2Wiki) (Ho et al., 2020), Musique (Trivedi et al., 2022b) and Bamboogle (Press et al., 2022). These datasets involve more complex multi-hop reasoning and are classified as out-of-domain because their question distributions differ significantly from our fine-tuning set. For testing, we randomly select 500 samples from each of the datasets, except for Bamboogle, where we use the entire 125 samples from its validation set.

### v3

@@ Abstract
Retrieval-Augmented Generation (RAG) has advanced open-domain question answering by incorporating external information into model reasoning. However, effectively leveraging external information to enhance reasoning presents the following challenges: (1) low signal-to-noise ratio, where answer-supportive external information is diluted by irrelevant material, and (2) error accumulation, which arises in multi-hop reasoning when incomplete or misleading information is incorporated. To address these challenges, we introduce EviNote-RAG, a framework that follows a retrieve–note–answer workflow. Instead of reasoning directly over raw external information, the model first produces Supportive-Evidence Notes (SENs), which concisely preserve answer-critical information and explicitly mark key and uncertainty information to improve accuracy. We further design an entailment-based Evidence Quality Reward (EQR) to ensure that SENs are logically sufficient to derive the final answer, thereby enhancing SENs’ quality. Experiments on both in-domain and out-of-domain QA benchmarks show that EviNote-RAG achieves state-of-the-art performance, improving answer accuracy, training stability, robustness, and efficiency. In particular, it yields relative F1 gains of 20% on HotpotQA (+0.093), 40% on Bamboogle (+0.151), and 91% on 2Wiki (+0.256), benefiting from improvements in the reasoning process.
@@ 1 Introduction
Large Language Models (LLMs) have evolved from next-token predictors into systems capable of advanced reasoning (Chowdhery et al., 2022; Verma et al., 2024b; Zheng et al., 2025; Yin et al., 2025; Jiang et al., 2025). Since the factual knowledge of LLMs is fixed during pre-training, they are prone to generating incorrect or outdated information when deployed in real-world tasks where knowledge evolves rapidly (Ji et al., 2022; Zhang et al., 2025). To address this limitation, Retrieval-Augmented Generation (RAG) (Arslan et al., 2024; Gao et al., 2025) has emerged by incorporating a search tool that supplies up-to-date external evidence at inference time, enabling models to ground their responses in timely information and improve factual consistency.
Despite recent advances, how to effectively leverage external documents to support reasoning remains a fundamental challenge (Gao et al., 2025). Prompt-based methods address this through multi-step reasoning (Jiang et al., 2024; Tran et al., 2024; Xiong et al., 2025) or adaptive workflows (Lee et al., 2024; Zhou et al., 2024; Wu et al., 2025a; Li et al., 2025a; Zhao et al., 2025b), while tuning-based strategies (Liu et al., 2024; Zhang et al., 2024a) improve fine-grained information extraction but often sacrifice generalization. More recently, advances in Reinforcement Learning (RL) (Kaelbling et al., 1996; Guo et al., 2025) have inspired RL-based RAG approaches (Zhang et al., 2024a; Wei et al., 2025a; Song et al., 2025b; Jin et al., 2025; Li et al., 2025a; Deng et al., 2025), which surpass earlier paradigms by exploring optimal strategies and enhancing generalization. Yet, RL-based RAG methods still rely on outcome-based rewards that evaluate only final correctness, offering little guidance for intermediate reasoning. Consequently, models remain constrained to the retrieve-then-answer paradigm, facing two persistent obstacles: (1) Low Signal-to-Noise Ratio (SNR), where retrieved evidence often includes substantial irrelevant content, making supportive information sparse (Shi et al., 2023; Jin et al., 2024); and (2) Error Accumulation, where reasoning errors (Shi et al., 2023) amplify when inference depends on incomplete or noisy evidence, especially in multi-hop QA. Addressing these issues calls for RL strategies that not only boost performance but also equip models with more effective workflows for handling long and noisy contexts.
@@
To address the core limitation of existing RAG systems, we propose EviNote-RAG, an end-to-end RL-based RAG framework that restructures the pipeline into a retrieve–note–answer process (Fig. 1). EviNote-RAG trains LLMs to generate Supportive-Evidence Notes (SENs), concise abstractions that preserve only answer-critical information and discard irrelevant content to improve answer accuracy. Each SEN further highlights key and uncertain information, echoing human note-taking strategies to improve focus and reduce misleading reasoning. Most importantly, we formulate evidence selection as a reinforcement learning problem: through the Evidence Quality Reward (EQR), a lightweight entailment judge evaluates how well each SEN supports the final answer. This reward signal encourages the model to explore strategies for evidence extraction while being guided toward more accurate and faithful use of information. As a result, EviNote-RAG achieves end-to-end optimization that reduces noise, mitigates error accumulation, and enables the model to learn an effective strategy for precise information utilization.
@@ 2 Related Work
2.1 Instruction-Guided RAG Methods
Instruction-guided methods (Amplayo et al., 2022; Yao et al., 2023; Jeong et al., 2024; Jiang et al., 2024; Wu et al., 2025a) enhance RAG by designing prompts that automate retrieval and guide multi-step reasoning (Jiang et al., 2024; Xiong et al., 2025). These approaches (Zhang et al., 2024b; Li et al., 2024a) typically decompose questions into sub-problems, retrieve external knowledge, and synthesize structured answers (Zhou et al., 2024; Zhao et al., 2025b; Tran et al., 2024). Other works (Li et al., 2025a; Lee et al., 2025) integrate retrieval directly into the reasoning loop, while more recent efforts (Yue et al., 2024; Verma et al., 2024a; Li et al., 2025a; Alzubi et al., 2025; Feng et al., 2025) interleave retrieval and reasoning adaptively. Despite these advances, prompt-based approaches inherently depend on the foundation model’s generalization ability in RAG, which remains limited. Our framework instead employs a post-training, reward-driven objective that explicitly shapes information selection and reasoning, leading to more faithful and task-adapted performance.
@@ 2.2 Reward-Guided Agentic RAG
Reward-guided approaches (Zhang et al., 2024a; Guan et al., 2025; Huang et al., 2025; Zhao et al., 2025a; Wang et al., 2024b; Wei et al., 2025a; Deng et al., 2025) employ Reinforcement Learning (RL) (Kaelbling et al., 1996) to optimize reasoning policies through scalar feedback derived from task performance. Early work (Nakano et al., 2021) demonstrated that reward signals can effectively guide multi-step retrieval and improve factual accuracy. Building on recent advances in RL (Guo et al., 2025), RL-based RAG approaches (Qi et al., 2025; Chen et al., 2025; Jin et al., 2025; Wei et al., 2025b) have surpassed previous paradigms by enabling optimal strategy exploration and improved generalization. Subsequent works have broadened this framework to diverse scenarios and tool use (Zheng et al., 2025; Dai et al., 2025; Wang et al., 2025; Sun et al., 2025a; Gutiérrez et al., 2025; Shao et al., 2025), highlighting the performance gains from RL-based supervision. However, most existing reward-guided methods (Wu et al., 2025b; Song et al., 2025b; Sun et al., 2025b) operate directly on raw, often noisy passages, which leads to a low signal-to-noise ratio (Shi et al., 2023; Jin et al., 2024) and error accumulation across multi-hop reasoning. EviNote-RAG tackles this limitation by using supportive-evidence notes to structure retrieved information and by applying supervision that enforces logical consistency between the notes and the final answers. These mechanisms together promote more reliable reasoning and improve answer accuracy.
@@ 3 Methodology
This section presents EviNote-RAG (as shown in Fig. 2), which integrates Supportive Evidence Notes (SENs) to distill answer-relevant content from retrievals and an Evidence Quality Reward (EQR) to ensure each note faithfully supports the final answer. Together, these components guide more accurate and robust reasoning. The following subsections describe the pipeline, SEN design, and reward formulation.
@@ 3.1 EviNote-RAG Pipeline
Upon receiving a query, EviNote-RAG either issues <search> to retrieve external evidence or, when sufficiently confident, answers directly. Retrieved content arrives as <information> and may be noisy; therefore, the system produces Supportive-Evidence Notes (SENs) in <summary> to filter distractors and retain evidence critical to the answer. Once sufficient notes are consolidated, the agent finalizes the response in <answer>. Importantly, SENs explicitly link supportive evidence to the evolving answer, ensuring consistency and precision. The following subsections provide further details.
3.2 Supportive-Evidence Note
To improve information utilization, our method uses Supportive-Evidence Notes (SENs) within <summary> tags to filter out irrelevant content, ensuring retention of supportive evidence. Next, we detail two key components of SENs: evidence-aware annotations and the dynamic SEN workflow, which jointly enhance information utilization and strengthen model performance.
Evidence-aware Annotations.
To enhance information utilization, SENs incorporate two annotation types: key information (denoted by * ) and uncertain information (denoted by – ). These annotations preserve the model’s certainty in multi-turn interactions, enabling precise identification of key information and avoiding misguidance from uncertain data, yielding substantial gains over naive summarization (Section. 4.6, Naive Summary vs. SEN).
Dynamic SEN Workflow.
We emphasize that SEN generation is optional after each retrieval phase, allowing the model to dynamically determine the necessity of summarization based on retrieval outcomes. This dynamic workflow design is essential for enhancing RL training effectiveness (Section. 4.6, Force Summary vs. Ours), underscoring the importance of flexibility in achieving optimal RAG strategies. Furthermore, To guide high-quality SEN generation, we propose the Evidence Quality Reward (EQR), which provides entailment-based feedback to focus on answer-relevant content. Details are presented in the next section.
3.3 Evidence Quality Reward
@@
To encourage the generation of high-quality SEN, the Evidence Quality Reward (EQR) introduces an Entailment Judge as a source of supervision. The underlying intuition is straightforward: a well-formed SEN should provide sufficient grounds for logically inferring the ground-truth answer. We realize the Entailment Judge through a lightweight Natural Language Inference (NLI) model (e.g., DistilBERT (Sanh et al., 2019)), which evaluates whether the final SEN entails the correct answer. To be more specific, we first construct an answer claim $h$ that asserts the ground-truth answer $\text{ANS}_{\text{gt}}$ is the correct answer to the question $q$. We then use the final SEN $\text{SEN}_{\text{last}}$ as the input text, and evaluate whether it logically supports $h$ using the Entailment Judge model $\mathcal{M}_{\text{Judge}}$, as shown below:
@@ 3.4 Training Strategy
Here, ✓ indicates satisfaction and ✗ a violation. The format criterion holds if the output includes an explicit <answer> tag, at least one <summary> tag, and follows the prescribed schema. The answer criterion requires an exact match with the ground truth, and the note-taking criterion holds when Evidence-aware Annotations are marked according to the SEN design. This reward design ensures that the model is gently guided to perform structured note-taking when its QA capability is still developing (through a small reward of $0.1+R_{\text{EQR}}$), and increasingly incentivized to generate precise, logically supported answers when it becomes more reliable (through $1+R_{\text{EQR}}$). Note that this strategy integrates the Evidence Quality Reward (EQR) $R_{\text{EQR}}$, which provides entailment-based feedback to further emphasize relevance and faithfulness in the final generated answers.
@@ 4.1 Datasets
We evaluated EviNote-RAG on seven widely used Question Answering (QA) benchmark datasets: (1) In-Domain Datasets consist of NQ (Kwiatkowski et al., 2019) and HotpotQA (Yang et al., 2018). These datasets are considered in-domain because they originate from the same question distribution, allowing for a direct comparison on familiar tasks. (2) Out-of-Domain Datasets: Out-of-Domain Datasets include PopQA (Mallen et al., 2022), TriviaQA (Joshi et al., 2017), 2WikiMultiHopQA (2Wiki) (Ho et al., 2020), Musique (Trivedi et al., 2022) and Bamboogle (Press et al., 2022). These datasets involve more complex multi-hop reasoning and are classified as out-of-domain because their question distributions differ significantly from our fine-tuning set. For testing, we randomly select 500 samples from each of the datasets, except for Bamboogle, where we use the entire 125 samples from its validation set.